In the configuration below, the battery is 50 V. Find the charge and energy stored on capacitor C2. 6. C3 C2
Answer #1

C2 and C3 are in parallel


C23 = C2 + C3

C23 = 7 uF


C23 and C1 are in series

1/Ceq = 1/C23 + 1/C1

1/Ceq = 1/7 + 1/2


Ceq = 1.56 uF


charge on C1 , Q1 = Ceq*V = 1.56*50 uC = 78 uC


potential across C2 , v2 = Q1/C23


charge on Capacitor C2 , Q2 = C2*V2 = C2*Q1/C23 = 3*78/7 uC = 33.4 uC


=========================


energy stored = (1/2)*Q2^2/C2 = (1/2)*(33.4*10^-6)^2/(3*10^-6) = 1.86*10^-4 J